# A Posterior Sampling Approach to Inferring Indoor Floorplans from Walk Trajectories

### v1

@@ 1 Introduction
Figure 2: (a) Fitting signal power using virtual $Tx$s. (b) Ideally, virtual $Tx$s should be located on wall surfaces. (c) Sparse virtual $Tx$s learnt by NeWRF shown in black/gray dots. (d) Dense virtual $Tx$s learnt by NeRF2. Neither correspond to the true (red) floorplan.
@@ 3 EchoNeRF Model
We train EchoNeRF on the measured data using our proposed objective function. This function only models the LoS and the first order reflections. We disregard the higher orders since they are very complex to model and contribute, on average, $<6\%$ of the total power (see statistics in Appendix C). The NeRF model we use is a remarkably simple MLP designed to predict the density $\delta\in[0,1]$ and orientation $\omega\in[-\pi,\pi]$ of a specified voxel in the indoor scene. The orientation aids in modeling reflections. The proposed objective function – parameterized by voxel attributes $\langle\delta,\omega\rangle$ and the $\langle Tx,Rx\rangle$ locations – models an approximation of the received power $\psi$ at that $Rx$ location. Minimizing $L_{2}$ loss of this power across all $Rx$ locations trains the MLP. Plotting out all the voxel densities in 2D gives us the estimated floorplan $F$.
@@ 3.4 Multi-stage Training
Stage 1: We first use the LoS model against the measured ground truth power $\psi^{*}$ (see Fig. 4). This converges quickly because the network easily learns the transparent voxels ($\delta$=00) located along LoS paths. For LoS paths that are occluded, the network incorrectly learns excessive opaque voxels between the $\langle Tx,Rx\rangle$, but this does not affect the LoS error since the path is anyway occluded. Hence, the outcome is a crude floorplan but a near-perfect LoS power estimate $\tilde{\psi}_{LoS}$. We utilize this $\tilde{\psi}_{LoS}$ in stage $2$ (discussed soon).
@@ 4 Experiments
Floor plan and Wireless Simulation Dataset. Floorplans are drawn from the Zillow Indoor Dataset  cruz2021zillow . In each floorplan, we use the $A^{*}$ algorithm  astar to generate a walking trajectory that traverses all rooms. We use the NVIDIA Sionna RT hoydis2022sionna ; hoydis2023sionna – a ray tracer for radio propagation modeling – to compute the ground truth signal power (also known as received signal strength index (RSSI)). We randomly place $M$ Txs, one in each room, denoted as $T_{m}$. To simulate omnidirectional transmissions at $2.4$GHz from each Tx location, we shoot $10^{7}$ rays into the given floorplan. For receiver locations, we sample the user trajectory at a fixed time interval to obtain $N$ Rx locations, denoted as $R_{n}$. Sionna accounts for specular reflections and refraction when these rays interact with walls in the specified floor plan; we use the default materials for the walls. As with most WiFi simulators, Sionna does not model signal penetration through walls – this means that a Tx and Rx located on opposite sides of a wall will not receive any RSSI. Overall, we gather $M\times N$ RSSI measurements $(T_{m},R_{n},\psi_{m,n})$ that serve as input to EchoNeRF.
@@ 4.2 Qualitative Results: Visual floorplans, RSSI heatmap, and basic ray tracing
Figure 6: Heatmaps highlighting EchoNeRF’s ability to learn signal propagation. (Top row) Inferred RSSI heatmaps with $Tx$ (red star) as used in training. (Bottom row) A new $Tx$ (green star) degrades NeRF2 and MLP while EchoNeRF shows accurate predictions.
@@ 4.3 Relaxing Assumptions & Sensitivity Study
$\blacksquare$ Receiver location error. Table 2 shows EchoNeRF’s sensitivity to $Rx$ location errors. We inject Gaussian noise $\mathcal{N}(0,\sigma^{2}I)$ to the $Rx$ locations; $\sigma=1$ implies a physical error of $1$m. Wall_IoU accuracy obviously drops with error but $0.5$ meter of error is tolerable without destroying the floorplan structure. Advancements in WiFi positioning systems have demonstrated robust sub-meter error.
$\blacksquare$ Effect of Furniture. Fig. 8 visualizes inferred floorplans when toy objects are scattered in open spaces ($Rx$ locations remain $N$=$2000$). EchoNeRF is able to identify some of the object blobs but sharpening the small objects is challenging due to more higher order reflections from furniture. Follow up work is needed, either in modeling $2^{nd}$ order reflections or by imposing stronger regularizations.
@@ Appendix B Modelling Wideband Multipath Signal Power
$=\operatorname{\mathbb{E}}\left[\sum_{l=0}^{L}|a_{lk}|^{2}\right]$
We used the fact that channel gains between different multipaths $l,m\ \text{with}\ l\neq m$ have vanishingly small correlation. Therefore, the total received power $\psi$ can be computed by summing all individual subcarrier powers.
@@ Appendix D Relaxing Tx Assumptions
We approximate that the $\psi_{i}$ is normally distributed with a mean modeled by the line-of-sight power $\frac{K}{d_{i}^{2}}$ and variance $\sigma^{2}$ where $d_{i}=||\mathbf{Tx}-\mathbf{Rx}_{i}||$ is the distance between $\mathbf{Tx}$ and $\mathbf{Rx}_{i}$. The likelihood function for each observation $\psi_{i}$ is thus given by:
@@ E.3 Detailed network parameters
We assume the floorplan is an unknown shape inside a $512\times 512$ grid. For any ray or ray segment, we uniformly sample $n_{r}=64$ voxels on it. We choose $\lambda_{1}=\lambda_{2}=0.01$ for LoS training followed by $\lambda_{1}=\lambda_{2}=0.1$ for training the EchoNeRF model. We find that discretized opacity values to $\{0,1\}$ improve our LoS model. We use the straight-through estimator [4] to avoid the unavailability of the gradient at the discretization step. To help optimization and to encourage sparsity of the number of reflections, we use only the top-$k$ contributions ($k$=$10$) while training the reflection model. We use the ADAM optimizer [17] with $1.0^{-4}$ learning rate. We train our models on NVIDIA A100 GPUs.

### v2

# Can NeRFs See without Cameras?

@@ 1 Introduction
Figure 2: (a) Fitting signal power using virtual TxTxs. (b) Ideally, virtual TxTxs should be located on wall surfaces. (c) Sparse virtual TxTxs learnt by NeWRF shown in black/gray dots. (d) Dense virtual TxTxs learnt by NeRF2. Neither correspond to the true (red) floorplan.
@@ 3 EchoNeRF Model
We train EchoNeRF on the measured data using our proposed objective function. This function only models the LoS and the first order reflections. We disregard the higher orders since they are very complex to model and contribute, on average, $6\%$ of the total power (see statistics in Appendix C). The NeRF model we use is a remarkably simple MLP designed to predict the density $\delta\in[0,1]$ and orientation $\omega\in[-\pi,\pi]$ of a specified voxel in the indoor scene. The orientation aids in modeling reflections. The proposed objective function – parameterized by voxel attributes $\langle\delta,\omega\rangle$ and the $\langle Tx,Rx\rangle$ locations – models an approximation of the received power $\psi$ at that $Rx$ location. Minimizing $L_{2}$ loss of this power across all $Rx$ locations trains the MLP. Plotting out all the voxel densities in 2D gives us the estimated floorplan $F$.
@@ 3.4 Multi-stage Training
Stage 1: We first use the LoS model against the measured ground truth power $\psi^{*}$ (see Fig. 4). This converges quickly because the network easily learns the transparent voxels ($\delta$=0) located along LoS paths. For LoS paths that are occluded, the network incorrectly learns excessive opaque voxels between the $\langle Tx,Rx\rangle$, but this does not affect the LoS error since the path is anyway occluded. Hence, the outcome is a crude floorplan but a near-perfect LoS power estimate $\tilde{\psi}_{LoS}$. We utilize this $\tilde{\psi}_{LoS}$ in stage $2$ (discussed soon).
@@ 4 Experiments
Floor plan and Wireless Simulation Dataset. Floorplans are drawn from the Zillow Indoor Dataset  cruz2021zillow . In each floorplan, we use the $A^{*}$ algorithm  astar to generate a walking trajectory that traverses all rooms. We use the NVIDIA Sionna RT hoydis2022sionna ; hoydis2023sionna – a ray tracer for radio propagation modeling – to compute the ground truth signal power (also known as received signal strength index (RSSI)). We randomly place $M$ Txs, one in each room, denoted as $T_{m}$. To simulate omnidirectional transmissions at 2.42.4GHz from each Tx location, we shoot $10^{7}$ rays into the given floorplan. For receiver locations, we sample the user trajectory at a fixed time interval to obtain $N$ Rx locations, denoted as $R_{n}$. Sionna accounts for specular reflections and refraction when these rays interact with walls in the specified floor plan; we use the default materials for the walls. As with most WiFi simulators, Sionna does not model signal penetration through walls – this means that a Tx and Rx located on opposite sides of a wall will not receive any RSSI. Overall, we gather $M\times N$ RSSI measurements $(T_{m},R_{n},\psi_{m,n})$ that serve as input to EchoNeRF.
@@ 4.2 Qualitative Results: Visual floorplans, RSSI heatmap, and basic ray tracing
Figure 6: Heatmaps highlighting EchoNeRF’s ability to learn signal propagation. (Top row) Inferred RSSI heatmaps with TxTx (red star) as used in training. (Bottom row) A new TxTx (green star) degrades NeRF2 and MLP while EchoNeRF shows accurate predictions.
@@ 4.3 Relaxing Assumptions & Sensitivity Study
$\blacksquare$ Receiver location error. Table 2 shows EchoNeRF’s sensitivity to $Rx$ location errors. We inject Gaussian noise $\mathcal{N}(0,\sigma^{2}I)$ to the $Rx$ locations; $\sigma=1$ implies a physical error of 11m. Wall_IoU accuracy obviously drops with error but $0.5$ meter of error is tolerable without destroying the floorplan structure. Advancements in WiFi positioning systems have demonstrated robust sub-meter error.
$\blacksquare$ Effect of Furniture. Fig. 8 visualizes inferred floorplans when toy objects are scattered in open spaces ($Rx$ locations remain NN=$2000$). EchoNeRF is able to identify some of the object blobs but sharpening the small objects is challenging due to more higher order reflections from furniture. Follow up work is needed, either in modeling $2^{nd}$ order reflections or by imposing stronger regularizations.
@@ Appendix B Modelling Wideband Multipath Signal Power
$=\operatorname{\mathbb{E}}\left[\sum_{l=0}^{L}|a_{lk}|^{2}\right]\$
[…]
@@ Appendix D Relaxing Tx Assumptions
[…]
@@ E.3 Detailed network parameters
We assume the floorplan is an unknown shape inside a $512\times 512$ grid. For any ray or ray segment, we uniformly sample $n_{r}=64$ voxels on it. We choose $\lambda_{1}=\lambda_{2}=0.01$ for LoS training followed by $\lambda_{1}=\lambda_{2}=0.1$ for training the EchoNeRF model. We find that discretized opacity values to $\{0,1\}$ improve our LoS model. We use the straight-through estimator [4] to avoid the unavailability of the gradient at the discretization step. To help optimization and to encourage sparsity of the number of reflections, we use only the top-$k$ contributions (kk=$10$) while training the reflection model. We use the ADAM optimizer [17] with $1.0^{-4}$ learning rate. We train our models on NVIDIA A100 GPUs.